### v1

# The associative-poset point of view on right regular bands

Abstract
We present two results on the relation between the class of right regular bands (RRBs) and their underlying associative posets. The first one is a construction of a left adjoint to the forgetful functor that takes an RRB $(P,\cdot)$ to the corresponding $(P,\leq)$. The construction of such a left adjoint is actually done in general for any class of relational structures $(X,R)$ obtained from a variety, where $R$ is defined by a finite conjunction of identities. The second result generalizes the “inner” representations of direct product decompositions of semilattices studied by the second author to RRBs having at least one commuting element.
footnotetext: Keywords: right posemigroup, adjoint functor theorem, factor congruence, inner direct product.footnotetext: MSC 2020: 06F05; 20M07.
1 Introduction
This paper is a continuation of our joint work with A. Petrovich [2], where we study right-regular bands (RRB) $(P,\cdot)$ using as a guiding concept the partial order which arise naturally from the stipulation
$x\leq y\iff x\cdot y=x.$ (1)
for all $x,y$ in $P$. The poset $(P,\leq)$ thus obtained is said to be associative.
We present two results on the relation between the class of RRBs and that of associative posets. The first one is motivated by the question on how to obtain a band from a given poset. To give some context, choosing an RRB structure for a given associative poset is not trivial, since this necessarily involves the Axiom of Choice as it was seen in [2, Thm. 6.2]. The answer we present here is to construct a left adjoint to the obvious forgetful functor that takes an RRB $(P,\cdot)$ to $(P,\leq)$ using (1). The construction of such a left adjoint is actually generalized to any class of relational structures $(X,R)$ defined from a variety using a finite conjunction of identities to define $R$. This is presented in Section 3, where we also discuss some other concrete cases of this construction (Section 3.1); notably, there is a corresponding adjunction between the category of RRBs and that of sets with an equivalence relation (Example 3.15).
The second result (Section 4) studies direct product representations of RRBs and generalizes [4] by the second author. The main motivation is to obtain “inner” representation of direct factors and factor congruences as they are generally obtained for some other algebraic structures as groups and rings (by way of direct summands). The reason to present this result here stems from the fact that the characterization of congruences and factors is informed by the underlying partial order. This was accomplished for RRBs having a non trivial “group-theoretic center;” that is, RRBs having at least one commuting element.
We conclude in Section 5 with some open questions.
2 Preliminaries
In this section we fix notation and introduce some basic notions.
We remind the reader that a right regular band (RRB) is an idempotent semigroup satifying
$a\cdot b\cdot a=b\cdot a.$ (2)
As mentioned in Section 1, a poset $(P,\leq)$ is said to be associative if there is an RRB operation such that (1) holds. We then use the phrase right posemigroup when referring to the expanded structure $(P,\leq,\cdot)$; this stems from the fact that the product is order-preserving on the right only (see Lemma 2.2.1). Conversely, given an RRB $(P,\cdot)$, we say that the partial order $x\leq y$ given by (1) is the underlying order of this RRB.
We state two simple lemmas whose proofs are left as an exercise for the interested reader, or can be checked at [2, Sect. 2].
Lemma 2.1.
In every right posemigroup,
  1. 1.
    $a\cdot b\leq b$; in particular, if $b$ is minimal, $a\cdot b=b$.
  2. 2.
    $a\leq b\implies b\cdot a=a$.∎
Lemma 2.2.
For every right posemigroup,
  1. 1.
    $a\leq b$ implies $a\cdot x\leq b\cdot x$.
  2. 2.
    $c\leq x,y$ implies $c\leq x\cdot y,y\cdot x$. Hence if $x\cdot y=y\cdot x$, it must be the infimum of $\{x,y\}$.∎
3 An adjunction
The categories of right-regular bands and associative posets are canonically linked by a forgetful functor. An adjunction between them can be constructed explicitly by making use of the algebraic structure of RRBs and associative posets. Instead of only doing this, in this section we present a more general result, by constructing left adjoints for a broad family of forgetful functors that includes the one mentioned above.
We fix a first-order language $L$ with an $n$-ary relation symbol $R$.
Definition 3.1.
Let $\mathcal{K}$ be a class of $L$-structures. We will say that $R$ is definable by conjunctions of identities over $\mathcal{K}$ if there exist $L$-terms $t_{1}(x_{1},\dots,x_{n})$, …, $t_{m}(x_{1},\dots,x_{n})$, $s_{1}(x_{1},\dots,x_{n})$, …, $s_{m}(x_{1},\dots,x_{n})$ such that
$\mathcal{K}\vDash\forall x_{1},\dots,x_{n},\,R(x_{1},\dots,x_{n})\iff\textstyle\bigwedge_{1\leq i\leq m}t_{i}(x_{1},\dots,x_{n})=s_{i}(x_{1},\dots,x_{n}).$
Definition 3.2.
Let $\mathcal{K}$ be a class of $L$-structures. We denote the set of function symbols of $L$ by $L_{\mathrm{alg}}$, and denote the reducts of $\mathcal{K}$ to $L_{\mathrm{alg}}$ and to $\{R\}$ by $\mathcal{K}_{\mathrm{alg}}$ and $\mathcal{K}_{R}$, respectively.
From now on, $\mathcal{K}$ will denote a class of $L$-structures such that $R$ is definable by conjunctions of identities over $\mathcal{K}$. Throughout this section we will study the interplay between the categories $\mathsf{K}_{\mathrm{alg}}$ and $\mathsf{K}_{R}$, which consist of the classes $\mathcal{K}_{\mathrm{alg}}$ and $\mathcal{K}_{R}$ and their homomorphisms, respectively.
Example 3.3.
Posemigroups provide an instance of Definition 3.1. Let $L:=\{\cdot,\leq\}$ and $\mathcal{K}$ be the class of right posemigroups. Here $R={\leq}$ is definable by one identity:
$\mathcal{K}\vDash\forall x_{1},x_{2},\ x_{1}\leq x_{2}\iff x_{1}\cdot x_{2}=x_{1}$
In this example, $\mathsf{K}_{\mathrm{alg}}$ is the variety of right regular bands and $\mathsf{K}_{R}$ is the category of associative posets.
There is a canonical forgetful functor $U:\mathsf{K}_{\mathrm{alg}}\rightarrow\mathsf{K}_{R}$ which assigns to every $\mathbf{A}\in\mathcal{K}_{\mathrm{alg}}$ the structure $\mathbf{A}_{R}:=(A,R^{\mathbf{A}})$ where
$\langle a_{1},\dots,a_{n}\rangle\in R^{\mathbf{A}}\iff\mathbf{A}\vDash\textstyle\bigwedge_{1\leq i\leq m}t_{i}(a_{1},\dots,a_{n})=s_{i}(a_{1},\dots,a_{n});$
and to every $\mathsf{K}_{\mathrm{alg}}$-morphism $f:\mathbf{A}\rightarrow\mathbf{B}$, the $\mathsf{K}_{R}$-morphism $Uf:\mathbf{A}_{R}\rightarrow\mathbf{B}_{R}$ given by $Uf(a)=f(a)$ (that is, $U$ preserves morphism as maps).
For the situation described in Example 3.3, the functor $U$ maps each RRB to its underlying associative poset.
We will show:
Theorem 3.4.
If $\mathcal{K}_{\mathrm{alg}}$ is a variety, the forgetful functor $U:\mathsf{K}_{\mathrm{alg}}\rightarrow\mathsf{K}_{R}$ has a left adjoint $F$.
We will start by defining the functor $F$ on objects.
For each set $X$, let’s consider the $\mathcal{K}_{\mathrm{alg}}$-free algebra over $X$, denoted by $\mathbf{F}_{\mathcal{K}}(X)$. We construe $\mathbf{F}_{\mathcal{K}}(X)$ as the quotient of the term algebra $T(X)$, by the least congruence $\delta_{X}$ such that the quotient belongs to $\mathcal{K}_{\mathrm{alg}}$. For $t\in T(X)$, we will denote by [[t]][\![t]\!] the equivalence class of $t$ under that congruence. In particular, $\{[\![x]\!]:x\in X\}$ is the set of free generators of $\mathbf{F}_{\mathcal{K}}(X)$.
We know that $\mathbf{F}_{\mathcal{K}}(X)$ has the following universal property: For every $\mathbf{B}\in\mathcal{K}_{\mathrm{alg}}$ and any function $\alpha:X\rightarrow B$, there exists a homomorphism $\beta:\mathbf{F}_{\mathcal{K}}(X)\rightarrow\mathbf{B}$ such that $\alpha(x)=\beta([\![x]\!])$ for every $x\in X$. We will use the following particular application:
Proposition 3.5.
Let $X$ and $Y$ be sets and $f:X\rightarrow Y$ a function. Then the map given by $\tilde{f}([\![t(\bar{x})]\!])=[\![t(f(\bar{x}))]\!]$ is a morphism $\tilde{f}:\mathbf{F}_{\mathcal{K}}(X)\rightarrow\mathbf{F}_{\mathcal{K}}(Y)$. ∎
Definition 3.6 ($F$ on objects).
For $\mathbf{X}\in\mathcal{K}_{R}$, let $\theta_{\mathbf{X}}$ be the congruence on $\mathbf{F}_{\mathcal{K}}(X)$ generated by
$\{\langle[\![t_{i}(x_{1},\dots,x_{n})]\!],[\![s_{i}(x_{1},\dots x_{n})]\!]\rangle\mid 1\leq i\leq m,\langle x_{1},\dots,x_{n}\rangle\in R^{\mathbf{X}}\}.$
The algebra $F\mathbf{X}$ is the quotient $\mathbf{F}_{\mathcal{K}}(X)/\theta_{\mathbf{X}}$.
We will also need Grätzer’s version of Mal’cev’s characterization of compact congruences.
Lemma 3.7.
Let $\mathbf{A}$ be any algebra and let $c,d\in A,$ $\bar{a},\bar{b}\in A^{j}.$ Then $\langle c,d\rangle\in\mathrm{Cg}^{A}(\bar{a},\bar{b})$ if and only if there exist $(j+m)$-ary terms $r_{1}(\bar{x},\bar{u}),\dots,r_{k}(\bar{x},\bar{u})$, with $k$ odd, and $\bar{z}\in A^{m}$ such that:
$c$ $=r_{1}(\bar{a},\bar{z})$
$r_{l}(\bar{b},\bar{z})$ $=r_{l+1}(\bar{b},\bar{z})$ $l\text{ odd,}$
$r_{l}(\bar{a},\bar{z})$ $=r_{l+1}(\bar{a},\bar{z})$ $l\text{ even,}$
$r_{k}(\bar{b},\bar{z})$ $=d.$
The following result completes the definition of $F$, now on arrows of $\mathsf{K}_{R}$.
Lemma 3.8 ($F$ on morphisms).
Let $\mathbf{X},\mathbf{Y}\in\mathcal{K}_{R}$ and $f:\mathbf{X}\rightarrow\mathbf{Y}$ an homomorphism. The function $Ff:F\mathbf{X}\rightarrow F\mathbf{Y}$ given by $Ff([\![t(\bar{x})]\!]/\theta_{\mathbf{X}})=[\![t(f(\bar{x}))]\!]/\theta_{\mathbf{Y}}$ is well defined and is a morphism.
Proof.
Let $f$ be as in the assumptions, and let $\tilde{f}:\mathbf{F}_{\mathcal{K}}(X)\rightarrow\mathbf{F}_{\mathcal{K}}(Y)$ be the morphism provided by Proposition 3.5 for $f$. In particular, we have $\tilde{f}([\![x]\!])=[\![f(x)]\!]$ for every $x\in X$. Now, take $c,d\in\mathbf{F}_{\mathcal{K}}(X)$ (where, $c=[\![t(\bar{x})]\!]$ and $d=[\![s(\bar{x})]\!]$ for some $s,t\in T(X)$) such that $\langle c,d\rangle\in\theta_{\mathbf{X}}$. If we can show that $\langle\tilde{f}(c),\tilde{f}(d)\rangle\in\theta_{\mathbf{Y}}$, we can then define $Ff([\![t(\bar{x})]\!]/\theta_{\mathbf{X}}):=\tilde{f}([\![t(\bar{x})]\!])/\theta_{\mathbf{Y}}$, and by how we have chosen $\tilde{f}$, this function will satisfy what we want. Since $\langle c,d\rangle\in\theta_{\mathbf{X}}$ is witnessed by finitely many tuples from the generators of $\theta_{\mathbf{X}}$, namely $\langle[\![t_{i_{p}}(x_{1}^{p},\dots,x_{n}^{p})]\!],[\![s_{i_{p}}(x_{1}^{p},\dots,x_{n}^{p})]\!]\rangle$ for some appropriate $x_{l}^{p}\in X$ (with $l=1,\dots,n$, $p=1,\dots,j$), Lemma 3.7 gives us $(j+m)$-ary terms $r_{1}(\bar{x},\bar{u})$,…, $r_{k}(\bar{x},\bar{u})$, with $k$ odd, and $\bar{z}\in\mathbf{F}_{\mathcal{K}}(X)^{m}$ such that:
$c$ $=r_{1}([\![\bar{W}]\!],\bar{z}),$ (3)
$r_{l}([\![\bar{V}]\!],\bar{z})$ $=r_{l+1}([\![\bar{V}]\!],\bar{z})$ for odd $l$, (4)
$r_{l}([\![\bar{W}]\!],\bar{z})$ $=r_{l+1}([\![\bar{W}]\!],\bar{z})$ for even $l$, (5)
$r_{k}([\![\bar{V}]\!],\bar{z})$ $=d$ (6)
where
$\bar{\mathrm{x}}$ $\mathrel{\mathop{:}}=x_{1}^{1},\dots,x_{n}^{1},\dots,x_{1}^{j},\dots,x_{n}^{j}$
$\bar{W}$ $=\bar{W}(\bar{\mathrm{x}})\mathrel{\mathop{:}}=t_{i_{1}}(x_{1}^{1},\dots,x_{n}^{1}),\dots,t_{i_{j}}(x_{1}^{j},\dots,x_{n}^{j})$
$\bar{V}$ $=\bar{V}(\bar{\mathrm{x}})\mathrel{\mathop{:}}=s_{i_{1}}(x_{1}^{1},\dots,x_{n}^{1}),\dots,s_{i_{j}}(x_{1}^{j},\dots,x_{n}^{j}),$
with the following abuses of notation: We take equivalence classes term-wise,
$\bar{W}]\!]=[\![t_{i_{1}}(x_{1}^{1},\dots,x_{n}^{1})]\!],\dots,[\![t_{i_{j}}(x_{1}^{j},\dots,x_{n}^{j})]\!],$
and apply functions in the same way:
$h(\bar{W})$ $=h\bigl(t_{i_{1}}(x_{1}^{1},\dots,x_{n}^{1})\bigr),\dots,h\bigl(t_{i_{j}}(x_{1}^{j},\dots,x_{n}^{j})\bigr)$
$\bar{W}\bigl(h(\bar{\mathrm{x}})\bigr)$ $=t_{i_{1}}\bigl(h(x_{1}^{1}),\dots,h(x_{n}^{1})\bigr),\dots,t_{i_{j}}\bigl(h(x_{1}^{j}),\dots,h(x_{n}^{j})\bigr)$
Using the fact that $\langle f(x_{1}^{i}),\dots,f(x_{n}^{i})\rangle\in R^{\mathbf{Y}}$ for every $i$, we will show that the terms obtained witness the fact that $\langle\tilde{f}(c),\tilde{f}(d)\rangle\in\theta_{\mathbf{Y}}$ by transitivity.
$\tilde{f}(c)$ $=\tilde{f}\bigl(r_{1}([\![\bar{W}(\bar{\mathrm{x}})]\!],\bar{z})\bigr)$ by (3)
$=r_{1}\bigl(\tilde{f}([\![\bar{W}(\bar{\mathrm{x}})]\!]),\tilde{f}(\bar{z})\bigr)$ $\tilde{f}$ is a morphism
$=r_{1}\bigl([\![\bar{W}(f(\bar{\mathrm{x}}))]\!],\tilde{f}(\bar{z})\bigr).$ by definition of $\tilde{f}$
Observe now that $\langle r_{1}([\![\bar{W}(f(\bar{\mathrm{x}}))]\!],\tilde{f}(\bar{z})),r_{1}([\![\bar{V}(f(\bar{\mathrm{x}}))]\!],\tilde{f}(\bar{z}))\rangle\in\theta_{\mathbf{Y}}$, where the second term appears for $l=1$ below. In general, for odd $l$,
$r_{l}\bigl([\![\bar{V}(f(\bar{\mathrm{x}}))]\!],\tilde{f}(\bar{z})\bigr)$ $=r_{l}\bigl(\bar{V}([\![f(\bar{\mathrm{x}})]\!]),\tilde{f}(\bar{z})\bigr)$ $\delta_{Y}$ is a congruence
$=r_{l}\bigl(\bar{V}(\tilde{f}([\![\bar{\mathrm{x}}]\!])),\tilde{f}(\bar{z})\bigr)$ by definition of $\tilde{f}$
$=r_{l}\bigl(\tilde{f}(\bar{V}([\![\bar{\mathrm{x}}]\!])),\tilde{f}(\bar{z})\bigr)$ $\tilde{f}$ is a morphism
$=r_{l}\bigl(\tilde{f}([\![\bar{V}(\bar{\mathrm{x}})]\!]),\tilde{f}(\bar{z})\bigr)$ $\delta_{X}$ is a congruence
$=\tilde{f}\bigl(r_{l}([\![\bar{V}(\bar{\mathrm{x}})]\!],\bar{z})\bigr)$ $\tilde{f}$ is a morphism
$=\tilde{f}\bigl(r_{l+1}([\![\bar{V}(\bar{\mathrm{x}})]\!],\bar{z})\bigr)$ by (4)
$=r_{l+1}\bigl(\tilde{f}([\![\bar{V}(\bar{\mathrm{x}})]\!]),\tilde{f}(\bar{z})\bigr)$ $\tilde{f}$ is a morphism
$=r_{l+1}\bigl(\tilde{f}(\bar{V}([\![\bar{\mathrm{x}}]\!])),\tilde{f}(\bar{z})\bigr)$ $\delta_{X}$ is a congruence
$=r_{l+1}\bigl(\bar{V}(\tilde{f}([\![\bar{\mathrm{x}}]\!])),\tilde{f}(\bar{z})\bigr)$ $\tilde{f}$ is a morphism
$=r_{l+1}\bigl([\![\bar{V}(f(\bar{\mathrm{x}}))]\!],\tilde{f}(\bar{z})\bigr)$ $\text{by definition of $\tilde{f}$}$.
In general, $\langle r_{l+1}([\![\bar{W}(f(\bar{\mathrm{x}}))]\!],\tilde{f}(\bar{z})),r_{l+1}([\![\bar{V}(f(\bar{\mathrm{x}}))]\!],\tilde{f}(\bar{z}))\rangle\in\theta_{\mathbf{Y}}$, for $l\leq k-1$. Similarly, for even $l$,
$r_{l}\bigl([\![\bar{W}(f(\bar{\mathrm{x}}))]\!],\tilde{f}(\bar{z})\bigr)$ $=r_{l}\bigl(\tilde{f}([\![\bar{W}(\bar{\mathrm{x}})]\!]),\tilde{f}(\bar{z})\bigr)$ by definition of $\tilde{f}$
$=\tilde{f}\bigl(r_{l}([\![\bar{W}(\bar{\mathrm{x}})]\!]),\bar{z}\bigr)$ $\tilde{f}$ is a morphism
$=\tilde{f}\bigl(r_{l+1}([\![\bar{W}(\bar{\mathrm{x}})]\!],\bar{z})\bigr)$ by (5)
$=r_{l+1}\bigl(\tilde{f}([\![\bar{W}(\bar{\mathrm{x}})]\!]),\tilde{f}(\bar{z})\bigr)$ $\tilde{f}$ is a morphism
$=r_{l+1}\bigl([\![\bar{W}(f(\bar{\mathrm{x}}))]\!],\tilde{f}(\bar{z})\bigr)$ $\text{by definition of $\tilde{f}$}$.
Finally,
$r_{k}\bigl([\![\bar{V}(f(\bar{\mathrm{x}}))]\!],\tilde{f}(\bar{z})\bigr)$ $=r_{k}\bigl(\bar{V}([\![f(\bar{\mathrm{x}})]\!]),\tilde{f}(\bar{z})\bigr)$ $\delta_{Y}$ is a congruence
$=r_{k}\bigl(\bar{V}(\tilde{f}([\![\bar{\mathrm{x}}]\!])),\tilde{f}(\bar{z})\bigr)$ by definition of $\tilde{f}$
$=r_{k}\bigl(\tilde{f}(\bar{V}([\![\bar{\mathrm{x}}]\!])),\tilde{f}(\bar{z})\bigr)$ $\tilde{f}$ is a morphism
$=r_{k}\bigl(\tilde{f}([\![\bar{V}(\bar{\mathrm{x}})]\!]),\tilde{f}(\bar{z})\bigr)$ $\delta_{X}$ is a congruence
$=\tilde{f}\bigl(r_{k}([\![\bar{V}(\bar{\mathrm{x}})]\!],\bar{z})\bigr)$ $\tilde{f}$ is a morphism
$=\tilde{f}(d)$ by (6).
Therefore the function $Ff$ is well defined. Now take $q$ an $n$-ary operation symbol in $L_{\mathrm{alg}}$. We have that
$Ff\bigl(q^{F\mathbf{X}}([\![u_{1}]\!]/\theta_{\mathbf{X}},\dots,[\![u_{n}]\!]/\theta_{\mathbf{X}})\bigr)$ $=\tilde{f}\bigl(q^{\mathbf{F}_{\mathcal{K}}(X)}([\![u_{1}]\!],\dots,[\![u_{n}]\!])/\theta_{\mathbf{X}}\bigr)$
$=q^{\mathbf{F}_{\mathcal{K}}(Y)}\bigl(\tilde{f}([\![u_{1}]\!]),\dots,\tilde{f}([\![u_{n}]\!])\bigr)/\theta_{\mathbf{Y}}$
$=q^{F\mathbf{Y}}\bigl(\tilde{f}([\![u_{1}]\!])/\theta_{\mathbf{Y}},\dots,\tilde{f}([\![u_{n}]\!])/\theta_{\mathbf{Y}}\bigr)$
$=q^{F\mathbf{Y}}\bigl(Ff([\![u_{1}]\!]/\theta_{\mathbf{X}}),\dots,Ff([\![u_{n}]\!]/\theta_{\mathbf{X}})\bigr)$
so $Ff$ is indeed a morphism. ∎
For a structure $\mathbf{X}\in\mathsf{K}_{R}$, we define an inclusion $\eta_{\mathbf{X}}:\mathbf{X}\to UF\mathbf{X}$ by $\eta_{\mathbf{X}}(a)\mathrel{\mathop{:}}=[\![a]\!]/\theta_{\mathbf{X}}$. The previous lemma allows us to prove the following:
Lemma 3.9.
For every $\mathbf{X}$ and $\mathbf{Y}$ in $\mathsf{K}_{R}$, and $\mathsf{K}_{R}$-morphism $h:\mathbf{Y}\rightarrow\mathbf{X}$:
$UFh\circ\eta_{\mathbf{Y}}=\eta_{\mathbf{X}}\circ h.$ (7)
Proof.
Let $y\in Y$. We have: $(\eta_{\mathbf{X}}\circ h)(y)=\eta_{\mathbf{X}}(h(y))=[\![h(y)]\!]/\theta_{\mathbf{X}}.$ Notice that, by the definition of $Fh$, this element is equal to $Fh([\![y]\!]/\theta_{\mathbf{Y}})=\bigl(Fh\circ\eta_{\mathbf{Y}}\bigr)(y)$. And since $U$ preserves morphisms as functions, we have that this is equal to $\bigl(UFh\circ\eta_{\mathbf{Y}}\bigr)(y)$. ∎
For a structure $\mathbf{X}\in\mathsf{K}_{R}$, we define a congruence $D_{\mathbf{X}}$ on $T(X)$ as follows: $\langle t(\bar{x}),s(\bar{x})\rangle\in D_{\mathbf{X}}$ if and only if $t^{\mathbf{A}}(\bar{x})=s^{\mathbf{A}}(\bar{x})$ for every $\mathbf{A}\in\mathcal{K}_{\mathrm{alg}}$ such that $U\mathbf{A}=\mathbf{X}$. It is clear that $\delta_{X}\subseteq D_{\mathbf{X}}$.
Lemma 3.10.
Given a structure $\mathbf{X}\in\mathsf{K}_{R}$, and an algebra $\mathbf{A}\in\mathcal{K}_{F}$ such that $U\mathbf{A}=\mathbf{X}$ there exists a morphism $g:F\mathbf{X}\rightarrow\mathbf{A}$ such that $g(\eta_{\mathbf{X}}(x))=x$ for every $x\in X$.
Proof.
Note there is a morphism $f:T(X)/D_{\mathbf{X}}\rightarrow\mathbf{A}$ for which $f(x/D_{\mathbf{X}})=x$, given by $f(t(\bar{x})/D_{\mathbf{X}})\mathrel{\mathop{:}}=t^{\mathbf{A}}(\bar{x})$; this is well defined by definition of $D_{\mathbf{X}}$.
We have that $\theta_{\mathbf{X}}\subseteq D_{\mathbf{X}}/\delta_{X}$: For every $\bar{x}\in X^{n}$ such that $\bar{x}\in R^{\mathbf{X}}$ we have that $\langle t_{i}(\bar{x}),s_{i}(\bar{x})\rangle\in D_{\mathbf{X}}$ for $1\leq i\leq m$ and hence
$\{\langle[\![t_{i}(\bar{x})]\!],[\![s_{i}(\bar{x})]\!]\rangle:1\leq i\leq m,\bar{x}\in R^{\mathbf{X}}\}\subseteq D_{\mathbf{X}}/\delta_{X}$
by definition of $D_{\mathbf{X}}/\delta_{X}$. Since $\theta_{\mathbf{X}}$ is generated by those pairs, we obtain the inclusion.
We now know that there is a projection morphism
$\pi:(T(X)/\delta_{X})/\theta_{\mathbf{X}}\rightarrow((T(X)/\delta_{X})/\theta_{\mathbf{X}})/((D_{\mathbf{X}}/\delta_{X})/\theta_{\mathbf{X}})$
which satisfies
$\pi([\![x]\!]/\theta_{\mathbf{X}})=([\![x]\!]/\theta_{\mathbf{X}})/\bigl((D_{\mathbf{X}}/\delta_{X})/\theta_{\mathbf{X}}\bigr).$
By the Second Isomorphism Theorem, there are isomorphisms
$h_{1}:\bigl((T(X)/\delta_{X})/\theta_{\mathbf{X}}\bigr)/\bigl((D_{\mathbf{X}}/\delta_{X})/\theta_{\mathbf{X}}\bigr)\rightarrow(T(X)/\delta_{X})/(D_{\mathbf{X}}/\delta_{X})$
and
$h_{2}:(T(X)/\delta_{X})/(D_{\mathbf{X}}/\delta_{X})\rightarrow T(X)/D_{\mathbf{X}}$
which satisfy
$h_{1}\bigl(([\![x]\!]/\theta_{\mathbf{X}})/\bigl((D_{\mathbf{X}}/\delta_{X})/\theta_{\mathbf{X}}\bigr)\bigr)=[\![x]\!]/(D_{\mathbf{X}}/\delta_{X})$
and
$h_{2}\bigl([\![x]\!]/(D_{\mathbf{X}}/\delta_{X})\bigr)=x/D_{\mathbf{X}}.$
We can now take $g\mathrel{\mathop{:}}=f\circ h_{2}\circ h_{1}\circ\pi:F\mathbf{X}\rightarrow A$ which is a morphism satisfying $g(\eta_{\mathbf{X}}(x))=x$. ∎
Lemma 3.11.
$F:\mathsf{K}_{R}\rightarrow\mathsf{K}_{\mathrm{alg}}$ is a functor.
Proof.
Given $\mathbf{X},\mathbf{Y},\mathbf{Z}\in\mathsf{K}_{R}$, and $f:\mathbf{X}\rightarrow\mathbf{Y}$, $g:\mathbf{Y}\rightarrow\mathbf{Z}$ $\mathsf{K}_{R}$-morphisms, we show that $Fg\circ Ff=F(g\circ f)$. Let $x\in F\mathbf{X}$ and assume $x=[\![t(\bar{x})]\!]/\theta_{\mathbf{X}}$ for some $t$ and $\bar{x}$. We have
$(Fg\circ Ff)(x)=Fg([\![t(f(\bar{x}))]\!]/\theta_{\mathbf{Y}})=t(gf(\bar{x}))/R_{Z})/\theta_{Z})=F(g\circ f)(x).$
Now let $\mathbf{X}\in\mathcal{K}_{R}$; we show that $F1_{\mathbf{X}}=1_{F\mathbf{X}}$. Take $x=[\![t(\bar{x})]\!]/\theta_{\mathbf{X}}\in F\mathbf{X}$. We have that
$F1_{\mathbf{X}}(x)=[\![t(1_{\mathbf{X}}(\bar{x}))]\!]/\theta_{\mathbf{X}}=[\![t(\bar{x})]\!]/\theta_{\mathbf{X}}=x.\qed$
Consider now the map $\varphi_{\mathbf{X},\mathbf{A}}:\mathsf{K}_{\mathrm{alg}}(F\mathbf{X},\mathbf{A})\to\mathsf{K}_{R}(\mathbf{X},U\mathbf{A})$ between homsets given by $\varphi_{\mathbf{X},\mathbf{A}}(f)=Uf\circ\eta_{\mathbf{X}}$.
Theorem 3.12.
The triple $\langle F,U,\varphi\rangle$ is an adjunction from $\mathsf{K}_{R}$ to $\mathsf{K}_{\mathrm{alg}}$.
Proof.
We first prove that $\varphi_{\mathbf{X},\mathbf{A}}$ is injective. Let $g:F\mathbf{X}\rightarrow\mathbf{A}$ and $h:F\mathbf{X}\rightarrow\mathbf{A}$ be two $\mathcal{K}_{F}$-morphisms such that $\varphi_{\mathbf{X},\mathbf{A}}(h)=\varphi_{\mathbf{X},\mathbf{A}}(g)$, i.e.,
$Uh(\eta_{\mathbf{X}}(x))=(Uh\circ\eta_{\mathbf{X}})(x)=(Ug\circ\eta_{\mathbf{X}})(x)=Ug(\eta_{\mathbf{X}}(x))$
for all $x\in\mathbf{X}$. From this we deduce that $g$ and $h$ coincide on all elements of $F\mathbf{X}$ of the form $x]\!]/\theta_{\mathbf{X}}=\eta_{\mathbf{X}}(x)$: We have that $h(\eta_{\mathbf{X}}(x))$ equals $Uh(\eta_{\mathbf{X}}(x))$ since $U$ preserves morphisms as functions. By hypothesis, $Uh(\eta_{\mathbf{X}}(x))=Ug(\eta_{\mathbf{X}}(x))$, the latter being equal to $g(\eta_{\mathbf{X}}(x))$ for the same reason. Therefore, $h(\eta_{\mathbf{X}}(x))=g(\eta_{\mathbf{X}}(x))$ for all $x\in\mathbf{X}$. Now, let $y\in F\mathbf{X}$ be of the form $x=[\![t(x_{1},\dots,x_{k})]\!]/\theta_{\mathbf{X}}$. Since $g$ and $h$ are morphisms, we have:
$g(y)$ $=g([\![t(x_{1},\dots,x_{k})]\!]/\theta_{\mathbf{X}})$
$=t^{\mathbf{A}}\bigl(g([\![x_{1}]\!]/\theta_{\mathbf{X}}),\dots,g([\![x_{k}]\!]/\theta_{\mathbf{X}})\bigr)$
$=t^{\mathbf{A}}\bigl(h([\![x_{1}]\!]/\theta_{\mathbf{X}}),\dots,h([\![x_{k}]\!]/\theta_{\mathbf{X}})\bigr)$
$=h([\![t(x_{1},\dots,x_{k})]\!]/\theta_{\mathbf{X}})$
$=h(y),$
which means that $g=h$.
Next we show that $\varphi_{\mathbf{X},\mathbf{A}}$ is surjective. Let $f:\mathbf{X}\rightarrow U\mathbf{A}$ be a $\mathsf{K}_{R}$-morphism. Take $Ff:F\mathbf{X}\rightarrow FU\mathbf{A}$ the $\mathsf{K}_{\mathrm{alg}}$-morphism given by Lemma 3.8, and compose it with the morphism $g:FU\mathbf{A}\rightarrow\mathbf{A}$ given by Lemma 3.10. Now we have $\varphi_{\mathbf{X},\mathbf{A}}(g\circ Ff)(x)=U(g\circ Ff)\circ\eta_{\mathbf{X}}(x)$. Since $U$ preserves morphisms as functions, this element is equal to $g(Ff(\eta_{\mathbf{X}}(x)))$ which, by the definition of $Ff$, is equal to $g(\eta_{U\mathbf{A}}(f(x)))$, and this last term is equal to $f(x)$ by construction of $g$. Therefore, $f=\varphi_{\mathbf{X},\mathbf{A}}(g\circ Ff)$.
Now let’s see that $\varphi$ satisfies the naturality conditions. That is, we want to show that for every pair of objects $\mathbf{X}\in\mathsf{K}_{R}$ and $\mathbf{A}\in\mathsf{K}_{\mathrm{alg}}$, every $\mathsf{K}_{\mathrm{alg}}$-morphism $f:F\mathbf{X}\rightarrow\mathbf{A}$, every $\mathsf{K}_{R}$-morphism $h:\mathbf{Y}\rightarrow\mathbf{X}$, and every $\mathsf{K}_{\mathrm{alg}}$-morphism $k:\mathbf{A}\rightarrow\mathbf{B}$, it holds that $\varphi_{\mathbf{X},\mathbf{B}}(k\circ f)=Uk\circ\varphi_{\mathbf{X},\mathbf{A}}(f)$ and $\varphi_{\mathbf{Y},\mathbf{A}}(f\circ Fh)=\varphi_{\mathbf{X},\mathbf{A}}f\circ h$.
On the one hand, we have $\varphi_{\mathbf{X},\mathbf{B}}(k\circ f)=U(k\circ f)\circ\eta_{\mathbf{X}}=Uk\circ Uf\circ\eta_{\mathbf{X}}=Uk\circ\varphi_{\mathbf{X},\mathbf{A}}(f)$, as desired.
On the other hand, we have $\varphi_{\mathbf{Y},\mathbf{A}}(f\circ Fh)=U(f\circ Fh)\circ\eta_{\mathbf{Y}}=Uf\circ UFh\circ\eta_{\mathbf{Y}}=Uf\circ\eta_{\mathbf{X}}\circ h=\varphi_{\mathbf{X},\mathbf{A}}(f)\circ h$, since $UFh\circ\eta_{\mathbf{Y}}=\eta_{\mathbf{X}}\circ h$ by Equation 7 from Lemma 3.9.
Therefore, the naturality conditions are satisfied. Thus, $\varphi$ is a natural bijection between $\mathsf{K}_{\mathrm{alg}}(F\mathbf{X},\mathbf{A})$ and $\mathsf{K}_{R}(\mathbf{X},U\mathbf{A})$, and hence the triple $\langle F,U,\varphi\rangle$ forms an adjunction. ∎
Corollary 3.13.
Let $\mathsf{RRB}$ and $\mathsf{AP}$ be the categories of RRBs with semigroup homomorphisms and associative posets with non decreasing maps, respectively. Then the forgetful functor $U:\mathsf{RRB}\to\mathsf{AP}$ admits a left adjoint.
Proof.
By Theorem 3.12 and Example 3.3. ∎
3.1 Examples
We now present some examples. In these examples, we omit the subscript in $\eta_{\mathbf{X}}$, as it will be clear from context.
Example 3.14.
In Figure 1 we can see an example of the partial order underlying $F\mathbf{P}$ for a simple poset $\mathbf{P}=(P,\leq)$. Each product of the two maximal elements in $UF\mathbf{P}$ is one of the two elements in the middle level (immediately below the second factor).
[…]
$\mathbf{P}$ $UF\mathbf{P}$
Figure 1:
Example 3.15.
In this example we show that an appropriate expansion $\mathcal{K}$ of the class of right-regular bands has $\mathcal{K}_{R}$ equal to the class of sets equipped with some equivalence relation.
Let $L:=\{\cdot,\theta\}$ and let $\mathcal{K}$ be the class satisfying all the right-regular band axioms together with the additional condition
$\forall x,y,\quad(x\mathrel{\theta}y)\Longleftrightarrow(x\cdot y=y)\wedge(y\cdot x=x).$
We have that $\mathcal{K}_{\text{alg}}$ is, once again, the class of right-regular bands. It can be proved that $\theta$ is always an equivalence relation.
Moreover, let $\delta$ be an equivalence relation over a set $S$. Index the equivalence classes of $\delta$ by a linearly ordered set $I$ and denote them by $\{x_{i}\mid i\in I\}$. Define an RRB operation $\cdot_{i}$ on each $x_{i}$ by
$a\cdot_{i}b:=b\quad\text{for all }a,b\in x_{i}.$
Now the ordered sum of the $x_{i}$ admits an RRB operation [2, Lemma 4.1]. It is clear that $a\cdot b=b\wedge b\cdot a=a$ holds if and only if $a,b\in x_{i}$ for some $i\in I$.
To illustrate this, we consider the three-element RRB $\mathbf{A}$ given by the multiplication table in Figure 2 (observe that the partial order underlying $\mathbf{A}$ is isomorphic to the poset $\mathbf{P}$ from the previous example). We have that $U\mathbf{A}$ is the set $\{a,b,0\}$ together with the equivalence relation given by the partition $\{\{a,b\},\{0\}\}$. We also provide the Hasse diagram describing the partial order underlying $FU\mathbf{A}$. Keep in mind that this partial order holds no relation to the original partial order.
$\cdot$ 0 $a$ $b$
0 0 0 0
$a$ 0 $a$ $b$
$b$ 0 $a$ $b$
$\cdot$ $a$ $b$ $c$ $w$ $x$ $y$ $z$
$a$ $a$ $b$ $y$ $w$ $x$ $y$ $z$
$b$ $a$ $b$ $z$ $w$ $x$ $y$ $z$
$c$ $w$ $x$ $c$ $w$ $x$ $y$ $z$
$w$ $w$ $x$ $y$ $w$ $x$ $y$ $z$
$x$ $w$ $x$ $z$ $w$ $x$ $y$ $z$
$y$ $w$ $x$ $y$ $w$ $x$ $y$ $z$
$z$ $w$ $x$ $z$ $w$ $x$ $y$ $z$
[…]
$\mathbf{A}$ $FU\mathbf{A}$ The partial order underlying $FU\mathbf{A}$
Figure 2:
Example 3.16.
Let $L:=\{\cdot,\leq\}$ and $\mathcal{K}$ the class of commutative right-posemigroups. $\mathcal{K}_{\text{alg}}$ is the class of meet-semilattices (as algebras) and $\mathcal{K}_{R}$ is the class of meet-semilattices (as posets).
Let $\mathbf{B}$ be the poset $\mathbf{P}$ in Example 3.14, regarded as a meet-semilattice. Then $U\mathbf{B}$ is that same poset and $FU\mathbf{B}$ is the four element meet-semilattice given by the Hasse diagram in Figure 3.
[…]
Figure 3: $FU\mathbf{B}$
Observe that $\mathbf{B}$ can also be thought of as an RRB, since semilattices are a subvariety of the variety of RRB’s. In that case, we would obtain the (different) $F\mathbf{P}$ from Example 3.14. This shows that the definition of the functor $F$ depends not only on the structure to which it is applied but also on the equational class.
Example 3.17.
Let $L:=\{\wedge,\vee,0,1,C\}$ and $K$ the class of bounded distributive lattices, along with the binary relation defined by “being a pair of complemented elements”. That is
$\mathcal{K}\models\forall x,y,C(x,y)\iff(x\wedge y=0)\;\&\;(x\vee y=1)$
Note that this relation is symmetric and irreflexive in every non-trivial lattice. Therefore, it can be represented as a graph. In this example, we consider the three-element chain as a lattice and show, in Figure 4, the graph corresponding to $U\mathbf{C}$ alongside the Hasse diagram of $FU\mathbf{C}$.
[…]
$\mathbf{C}$ $U\mathbf{C}$ $FU\mathbf{C}$
Figure 4:
Observe that any graph obtained in this way necessarily includes a two-vertex component (namely, the one formed by the vertices corresponding to 0 and $1$).
3.2 SAFT applicability
Although one might expect a result such as Corollary 3.13 to follow from general categorical principles (such as the Special Adjoint Functor Theorem [3, Sect. V.8]), this is not the case as we now show. Specifically, the adjunction we construct cannot be derived from that theorem, as the category of RRBs lacks a cogenerating set, a situation common to residually large idempotent varieties. For completeness, we include a proof of this last fact.
Lemma 3.18.
Let $\mathcal{V}$ a idempotent variety. If $\mathcal{V}$ has a cogenerating set, then it is residually small.
Proof.
Suppose there exists a cogenerating set $Q$. Let $\mathbf{B}\in\mathcal{V}$. Let $\operatorname{\mathrm{Hom}}(\mathbf{B},\mathbf{Q})$ be the set of homomorphisms from $\mathbf{B}$ to $\mathbf{Q}$. Define $\mathbf{P}$ as
$\mathop{\mathbf{P}:=\prod_{\mathbf{Q}\in Q}\prod_{g\in\operatorname{\mathrm{Hom}}(\mathbf{B},\mathbf{Q})}\mathbf{Q}}$
and let $f:\mathbf{B}\rightarrow\mathbf{P}$ be defined coordinate-wise as $f(x)_{\mathbf{Q},g}=g(x)$. We claim that $f$ is an embedding from $\mathbf{B}$ into $\mathbf{P}$. Indeed, let $x,y\in\mathbf{B}$ with $x\neq y$. Let $\{*\}$ be the trivial algebra and define $h,h^{\prime}:\{*\}\rightarrow\mathbf{B}$ as the homomorphisms with image $\{x\}$ and $\{y\}$ respectively. Since $Q$ is a cogenerating set, there exist $\mathbf{Q}\in Q$ and $g:\mathbf{B}\rightarrow\mathbf{Q}$ such that $g\circ h\neq g\circ h^{\prime}$ hence $g(x)\neq g(y)$. Therefore, we have that $f(x)_{\mathbf{Q},g}\neq f(y)_{\mathbf{Q},g}$ and hence $f(x)\neq f(y)$.
By Birkhoff’s Theorem there exists a suitable set and a subdirect embedding $\alpha_{\mathbf{Q}}:\mathbf{Q}\rightarrow\prod_{i\in I_{\mathbf{Q}}}\mathbf{A}_{i}$. Now, we can define a new subdirect embedding
$\mathop{\alpha:\mathbf{P}\longrightarrow\prod_{\mathbf{Q}\in Q}\prod_{g\in\operatorname{\mathrm{Hom}}(\mathbf{B},\mathbf{Q})}\prod_{i\in I_{\mathbf{Q}}}\mathbf{A}_{i}}$
given by $\alpha(x)_{\mathbf{Q},g,i}=\alpha_{\mathbf{Q}}(x_{\mathbf{Q},g})_{i}$. Define $\mathcal{S}_{\mathbf{Q},g,i}$ to be the image of $\pi_{\mathbf{Q},g,i}\circ\alpha\circ f$ where
$\mathop{\pi_{\mathbf{Q},g,i}:\prod_{\mathbf{Q}\in Q}\prod_{g\in\operatorname{\mathrm{Hom}}(\mathbf{B},\mathbf{Q})}\prod_{j\in I_{\mathbf{Q}}}\mathbf{A}_{j}}\longrightarrow\mathbf{A}_{i}$
is the canonical projection for all $\mathbf{Q}\in Q$, $g\in\operatorname{\mathrm{Hom}}(\mathbf{B},\mathbf{Q})$, $i\in I_{\mathbf{Q}}$. Note that each $\mathcal{S}_{\mathbf{Q},g,i}$ is a substructure of $\mathbf{A}_{i}$ and that the function
$\mathop{\beta:b\longrightarrow\prod_{\mathbf{Q}\in Q}\prod_{g\in\operatorname{\mathrm{Hom}}(\mathbf{B},\mathbf{Q})}\prod_{i\in I_{\mathbf{Q}}}}\mathcal{S}_{\mathbf{Q},g,i}$
given by $\beta(x)_{\mathbf{Q},g,i}=(\pi_{\mathbf{Q},g,i}\circ\alpha\circ f)(x)$ is a subdirect embedding.
By the previous discussion, we have that each $\mathbf{B}\in\mathcal{V}$ is isomorphic to a subdirect product of algebras in $\mathcal{S}(\{\mathbf{A}_{i}\mid\mathbf{Q}\in Q,i\in I_{\mathbf{Q}}\})$ where this last family is the set of substructures of some $\mathbf{A}_{i}$ with $\mathbf{Q}\in Q$ and $i\in I_{\mathbf{Q}}$. Therefore, every subdirectly irreducible algebra must be isomorphic to some member in $\mathcal{S}(\{\mathbf{A}_{i}\mid\mathbf{Q}\in Q,i\in I_{\mathbf{Q}}\})$. This shows that $\mathcal{V}$ must be residually small. ∎
Lemma 3.19.
The variety of RRB’s is residually large.
The proof presented here follows an example given in [1]. To proceed, we will rely on an auxiliary result, whose proof is provided in the cited source. We begin by defining two fundamental concepts.
Definition 3.20.
  • A zero of a semigroup $\mathbf{S}$ is an element $0\in\mathbf{S}$ which satisfies $0\cdot s=s\cdot 0=0$ for all $s\in\mathbf{S}$
  • Given a semigroup $\mathbf{S}=(S,\cdot)$, its dual $\mathbf{S}^{*}$ is the semigroup with multiplication given by $x\tilde{\cdot}y=y\cdot x$ for all $x,y\in S$.
Theorem 3.21.
An idempotent semigroup $S$ (without zero) is subdirectly irreducible if and only if $\mathbf{S}$ or its dual $\mathbf{S}^{*}$ is isomorphic to a semigroup $\mathbf{T}$ which satisfies the following two conditions:
  • $C(X)\subseteq\mathbf{T}\subseteq X^{X}$, where $X^{X}$ is the semigroup of all mappings of $X$ into itself, $C(X)$ is the set of all constant mappings of $X$, and $\mathbf{T}$ is a subsemigroup of $X^{X}.$
  • There exist $k,k^{\prime}\in C(X)$ such that $\theta(k,k^{\prime})\subseteq\theta(c,d)$ for all $\in C(X), cdc\neq d$, where $\theta(x,y)$ denotes the smallest congruence containing (x,y)(x,y).∎
Proof of Lemma 3.19.
We will show that subdirectly irreducible right-regular bands of arbitrarily large cardinality exist. Fix an arbitrary set $X$ with two distinguished elements 0 and $1$. Let $A$ consist of the following functions from $X$ into itself:
  • The mappings $\{a_{i}\mid i\in X\}$ are the constant mappings defined by $a_{i}(j)=i$.
  • The mappings $\{b_{i}\mid i\in X,i\notin\{0,1\}\}$ are defined by $b_{i}(i)=i$, $b_{i}(j)=0$ if $i\neq j$.
  • The mappings $\{c_{i}\mid i\in X,i\notin\{0,1\}\}$ are defined by $c_{i}(i)=i$, $c_{i}(j)=1$ if $i\neq j$.
Define multiplication on $A$ by $a\cdot b=b\circ a$ where $\circ$ denotes composition of functions. Since the $a_{i},b_{i}$, and $c_{i}$ are idempotent, we can show that it is an idempotent semigroup by proving that it is closed with respect to the composition of functions. If one of the factors in a given composition is a constant, the result is a constant. If both are non-constants, they occur among the following: $c_{i}\circ b_{i}=c_{i}$, $b_{i}\circ c_{i}=b_{i}$, and if $i\neq j$, $c_{j}\circ b_{i}=c_{j}\circ c_{i}=a_{1}$, $b_{j}\circ b_{i}=b_{j}\circ c_{i}=a_{0}$. We will now consider the dual $\mathbf{A}^{*}$, where multiplication is given by composition of functions. We will show that this semigroup satisfies the conditions of Theorem 3.21, which tells us that $\mathbf{A}$ is subdirectly irreducible (note that since $a_{i}\cdot a_{j}=a_{j}$ for all $i,j$, $\mathbf{A}$ has no zero). By the mentioned theorem, it is enough to show that $\theta(a_{0},a_{1})\subseteq\theta(a_{i},a_{j})$ for all $i\neq j$. If $i\neq j$, $\theta(a_{i},a_{j})$ contains the pairs $(b_{i}\circ a_{i},b_{i}\circ a_{j})=(a_{i},a_{0})$ and $(c_{i}\circ a_{i},c_{i}\circ a_{j})=(a_{i},a_{0})$, and therefore it must contain $(a_{0},a_{1})$, thus $\mathbf{A}$ is a subdirectly irreducible idempotent semigroup. We now show that $\mathbf{A}$ satisfies $x\cdot y\cdot x=y\cdot x)$. If $x,y$ or $x\cdot y$ is a constant, or if $x=y$, the result is trivial. The only other case is $\{x,y\}=\{b_{i},c_{j}\}$ for some $i,j$ and in that case it is easy to check that $x\cdot y\cdot x=y\cdot x$. ∎
We conclude that there’s no cogenerating set for the variety of RRB’s, so the Special Adjoint Functor Theorem cannot be applied in this case.
4 Factor congruences of right-regular bands
In the following, we provide an application of the order-theoretic perspective on bands to the study of direct product decompositions of RRBs with a central element. This generalizes the results from [4].
One key idea from that paper (which studies join-semilattices) was that existing binary infima in finite direct products of join-semilattices must factorize and satisfy some distributive and absorption properties with respect to the join. Since the $(\leq,\cdot)$-posemigroups associated to commutative bands correspond to meet-semilattices (Lemma 2.2(2)), all concepts here will be dual those in [4]; in particular, we will be speaking of partial binary suprema. We have the analogous
Lemma 4.1.
For any pair of RRBs $C,D$ and elements $c,e\in C$ and $d,f\in D$, the element $\langle c,d\rangle\vee\langle e,f\rangle$ exists in $C\times D$ if and only if $c\vee e$ and $d\vee f$ exist.
The main difference between this section and [4] is that, unlike $\vee$ nor $\wedge$, the RRB operation $\cdot$ is not commutative in general. We do assume for the rest of this section that at least one element commutes with every other:
Definition 4.2.
Given an RRB $(A,\cdot)$, an element $c\in A$ is said to be central if for all $x\in A$, we have $x\cdot c=c\cdot x$.
We will highlight modifications to the proofs in the sequel.
4.1 Decomposition into direct product of RRBs
For the remainder of this chapter, we will write formulas in the language $\{\cdot,\vee\}$; occurrences of $\leq$ can be reduced using (1). The formula $x\vee y=z$” will be interpreted as “$z$ is the supremum of $\{x,y\}$”:
$x,y\leq z\mathrel{\&}(\forall u:x,y\leq u\;\rightarrow\;z\leq u),$
unless otherwise specified; in general, every equation $t_{1}=t_{2}$ involving $\vee$ should be interpreted as “if one of the terms exists, so does the other and they are equal.” It is easy to check that the laws of associativity hold for the partial operation $\vee$ in every poset.
From now on, let $A$ be an RRB and $c\in A$ be a fixed central element. Let $\varphi(c,x_{1},x_{2},x)$ be the conjunction of the following formulas:
  • comm
    $x\cdot x_{1}=x_{1}\cdot x$ and $x\cdot x_{2}=x_{2}\cdot x$.
  • dist
    $x=(x\cdot x_{1})\vee(x\cdot x_{2}).$
  • p1
    $x_{1}=(x\cdot x_{1})\vee(c\cdot x_{1}).$
  • p2
    $x_{2}=(x\cdot x_{2})\vee(c\cdot x_{2}).$
  • prod
    $x_{1}\cdot x_{2}=x\cdot c.$
Note that comm is the only formula that does not appear in the formula $\varphi(c,x_{1},x_{2},x)$ defined in [4]. Note also that, due to the assumption that $c$ is central and comm, Lemma 2.2(2) guarantees that all products appearing in $\varphi$ except for the last one are actually infima. We will write $x=\langle\langle x_{1},x_{2}\rangle\rangle_{c}$” to denote that $\varphi(c,x_{1},x_{2},x)$ holds.
Definition 4.3.
Suppose that $I_{1},I_{2}$ are subsemigroups of $A$. We will say that $A$ is the $c$-direct product of $I_{1}$ and $I_{2}$, and write $A=I_{1}\times_{c}I_{2}$, if and only if the following conditions hold:
  • Perm
    The elements of $I_{1}$ commute with those of $I_{2}$.
  • Mod1
    For all $x,y\in A$, $x_{1}\in I_{1}$, and $x_{2}\in I_{2}$, if $x\cdot c\leq x_{1}\cdot x_{2}$, then
    $\bigl((x\cdot x_{1})\vee(x\cdot x_{2})\bigr)\cdot y=(x\cdot x_{1}\cdot y)\vee(x\cdot x_{2}\cdot y),$
    $y\cdot\bigl((x\cdot x_{1})\vee(x\cdot x_{2})\bigr)=(y\cdot x\cdot x_{1})\vee(y\cdot x\cdot x_{2}).$
  • Mod2
[…]
    $\bigl((x\cdot x_{i})\vee(c\cdot x_{i})\bigr)\cdot y=(x\cdot x_{i}\cdot y)\vee(c\cdot x_{i}\cdot y),$
    $y\cdot\bigl((x\cdot x_{i})\vee(c\cdot x_{i})\bigr)=(y\cdot x\cdot x_{i})\vee(y\cdot c\cdot x_{i}).$
    for $i=1,2$.
  • Abs
    For all $x_{1},y_{1}\in I_{1}$, and $z_{2}\in I_{2}$, we have: $x_{1}\vee(y_{1}\cdot z_{2})=x_{1}\vee(y_{1}\cdot c)$ (and exchanging the roles of $I_{1}$ and $I_{2}$).
  • Exi
    $\forall x_{1}\in I_{1},x_{2}\in I_{2}\;\exists x\in A:x=\langle\langle x_{1},x_{2}\rangle\rangle_{c}.$
  • Onto
    $\forall x\in A\ \exists x_{1}\in I_{1},x_{2}\in I_{2}:\;x=\langle\langle x_{1},x_{2}\rangle\rangle_{c}.$
Note that these are “two-sided” versions of the conditions in [4], plus the commutativities Perm.
For the sake of brevity, we will omit the reference to $c$, writing $x=\langle\langle x_{1},x_{2}\rangle\rangle$ instead of $x=\langle\langle x_{1},x_{2}\rangle\rangle_{c}$. Note that Exi implies:
  • Ori
    $\forall x_{1}\in I_{1},x_{2}\in I_{2}\;(x_{1}\cdot x_{2}\leq c)$.
4.2 The representation lemma
Lemma 4.4.
Suppose $A=I_{1}\times_{c}I_{2}$. Then $x=\langle\langle x_{1},x_{2}\rangle\rangle$ defines an isomorphism $\langle x_{1},x_{2}\rangle\stackrel{{\scriptstyle\varphi}}{{\mapsto}}x$ between $I_{1}\times I_{2}$ and $A$.
Proof.
The proof that the function $\langle x_{1},x_{2}\rangle\stackrel{{\scriptstyle\varphi}}{{\mapsto}}x$ is well-defined is very similar to that in [4], but it involves essential uses of Perm and comm. It is left as an interesting exercise for the reader.
The function defined by $\varphi$ is surjective by Onto; to see it is injective, let $x=\langle\langle x_{1},x_{2}\rangle\rangle$ and $x=\langle\langle y_{1},y_{2}\rangle\rangle$. We have:
$x_{1}$ $=x_{1}\vee(x_{1}\cdot x_{2})$
$=x_{1}\vee(x\cdot c)$ by prod
$=x_{1}\vee(y_{1}\cdot y_{2})$ by prod again
$=x_{1}\vee(y_{1}\cdot c)$ by Abs
and then $x_{1}\geq y_{1}\cdot c$. Similarly, $y_{1}\geq x_{1}\cdot c$ and in conclusion, by Lemma  2.2(1),
$x_{1}\cdot c=y_{1}\cdot c.$ (8)
On the other hand,
$x\cdot y_{1}$ $=y_{1}\cdot x$ by comm
$=y_{1}\cdot((x\cdot x_{1})\vee(x\cdot x_{2}))$ by dist
$=(y_{1}\cdot x\cdot x_{1})\vee(y_{1}\cdot x\cdot x_{2})$ by Mod1
$=(x\cdot y_{1}\cdot x_{1})\vee(x\cdot y_{1}\cdot x_{2})$ by comm
$=(x\cdot y_{1}\cdot x_{1})\vee(x\cdot y_{1}\cdot x_{2}\cdot c)$ by Ori
$=(x\cdot y_{1}\cdot x_{1})\vee(x\cdot c)$ $c$ central and $x\cdot c\leq x_{1},y_{1},x_{2},y_{2}$
$=x\cdot y_{1}\cdot x_{1},$ by the same argument.
and also
$x\cdot x_{1}$ $=((x\cdot y_{1})\vee(x\cdot y_{2}))\cdot x_{1}$ by dist
$=(x\cdot y_{1}\cdot x_{1})\vee(x\cdot y_{2}\cdot x_{1})$ by Mod1
$=(x\cdot y_{1}\cdot x_{1})\vee(x\cdot x_{1}\cdot y_{2})$ by Perm
$=(x\cdot y_{1}\cdot x_{1})\vee(x\cdot x_{1}\cdot y_{2}\cdot c)$ by Ori
$=(x\cdot y_{1}\cdot x_{1})\vee(x\cdot c)$ $c$ central and $x\cdot c\leq x_{1},y_{1},x_{2},y_{2}$
$=x\cdot y_{1}\cdot x_{1}.$ by the same argument.
Thus, we obtain $x\cdot x_{1}=x\cdot y_{1}$. (Here, we must use the fact that $x\cdot y_{1}=y_{1}\cdot x$ because otherwise, it is not possible to prove that $x\cdot y_{1}=x\cdot x_{1}$). Combining this with (8) and using p1, we have
$x_{1}=(x\cdot x_{1})\vee(c\cdot x_{1})=(x\cdot y_{1})\vee(c\cdot y_{1})=y_{1}.$
By the same reasoning, $x_{2}=y_{2}$. The preceding part requires a more extensive development than the proof in [4].
Next we prove that $\varphi$ preserves $\cdot$. Suppose $x=\langle\langle x_{1},x_{2}\rangle\rangle$ and $z=\langle\langle z_{1},z_{2}\rangle\rangle$; since each $I_{1},I_{2}$ is a subsemigroup, we know that $x_{j}\cdot z_{j}\in I_{j}$ for $j=1,2$. We want to show $x\cdot z=\langle\langle x_{1}\cdot z_{1},x_{2}\cdot z_{2}\rangle\rangle$. The property prod is immediate (and therefore, we can apply Mod1 and Mod2 to $x\cdot z$). Now we prove dist:
$x\cdot z$ $=\bigl((x\cdot x_{1})\vee(x\cdot x_{2})\bigr)\cdot z$ by dist for $x$
$=(x\cdot x_{1}\cdot z)\vee(x\cdot x_{2}\cdot z).$ by Mod1
This last term is equal to
$(x\cdot x_{1}\cdot z\cdot z_{1})\vee(x\cdot x_{1}\cdot z\cdot z_{2})\vee(x\cdot x_{2}\cdot z\cdot z_{1})\vee(x\cdot x_{2}\cdot z\cdot z_{2}),$ (9)
by dist for $z$ and Mod1. Note that, for any $y$,
$(y\cdot z\cdot z_{1})$ $=y\cdot z\cdot(z\cdot z_{1}\vee c\cdot z_{1})$ by p1 for $z_{1}$
$=(y\cdot z\cdot z_{1})\vee(y\cdot z\cdot c\cdot z_{1})$ by Mod1 for $z$
$=(y\cdot z\cdot z_{1})\vee(y\cdot z_{1}\cdot z_{2}\cdot z_{1})$ by prod for $z$
$=(y\cdot z\cdot z_{1})\vee(y\cdot z_{2}\cdot z_{1})$ by Eq. 2
$=(y\cdot z\cdot z_{1})\vee(z_{2}\cdot y\cdot z_{1})$ by Perm
$=(y\cdot z\cdot z_{1})\vee(z_{1}\cdot z_{2}\cdot y\cdot z_{1})$ by Eq. 2 again
$=(y\cdot z\cdot z_{1})\vee(z\cdot c\cdot y\cdot z_{1})$ by prod for $z$
$=(z\cdot y\cdot z\cdot z_{1})\vee(z\cdot y\cdot c\cdot z_{1})$ by Eq. 2; $c$ is central
$=z\cdot y\cdot(z\cdot z_{1}\vee c\cdot z_{1})$ by Mod1 for $z$
$=z\cdot y\cdot z_{1}.$
This development is necessary in order to permute $x_{1}$ and $x_{2}$ with $z$ in the first and fourth term of (9) respectively, as we cannot guarantee that $\cdot$ commutes over these elements. This equality will be used multiple times in this proof, always for the same reason. Similarly, $z\cdot y\cdot z_{2}=y\cdot z\cdot z_{2}$.
We can now rewrite the term (9) as follows:
$(x\cdot z\cdot x_{1}\cdot z_{1})\vee(x\cdot x_{1}\cdot z\cdot z_{2})\vee(x\cdot x_{2}\cdot z\cdot z_{1})\vee(x\cdot z\cdot x_{2}\cdot z_{2}).$
Note that
$x\cdot x_{1}\cdot z\cdot z_{2}$ $=x\cdot x_{1}\cdot z_{2}\cdot z$ by comm
$=x\cdot x_{1}\cdot z_{2}\cdot c\cdot z$ by Ori
$=x\cdot c\cdot x_{1}\cdot z_{2}\cdot z_{1}\cdot z_{2}$ by Eq. 2 and $c$ is central
$=x_{1}\cdot x_{2}\cdot x_{1}\cdot z_{2}\cdot z_{1}\cdot z_{2}$ by prod
$=x_{2}\cdot x_{1}\cdot z_{1}\cdot z_{2}$
$=x_{1}\cdot x_{2}\cdot z_{1}\cdot z_{2}.$ by Perm
Similarly, $x\cdot x_{2}\cdot z\cdot z_{1}=x_{1}\cdot x_{2}\cdot z_{1}\cdot z_{2}$
Let’s see that $(x_{1}\cdot x_{2}\cdot z_{1}\cdot z_{2})\leq(x\cdot x_{1}\cdot z\cdot z_{1})$:
$(x_{1}\cdot x_{2}\cdot z_{1}\cdot z_{2})\cdot(x\cdot z\cdot x_{1}\cdot z_{1})$ $=x_{2}\cdot z_{2}\cdot x\cdot x_{1}\cdot z\cdot z_{1}$
$=x_{1}\cdot x_{2}\cdot z_{2}\cdot x\cdot z\cdot z_{1}$ by comm, Perm
$=x\cdot c\cdot z_{2}\cdot x\cdot z\cdot z_{1}$ by prod for $x$
$=z_{2}\cdot x\cdot z\cdot c\cdot z_{1}$
$=z_{2}\cdot x\cdot c\cdot z\cdot c\cdot z_{1}$
$=z_{2}\cdot x_{1}\cdot x_{2}\cdot z_{1}\cdot z_{2}\cdot z_{1}$ by prod for $z$
$=x_{1}\cdot x_{2}\cdot z_{2}\cdot z_{1}$
$=x_{1}\cdot x_{2}\cdot z_{1}\cdot z_{2}.$
Therefore, we can eliminate the two middle terms in Equation (9) and obtain dist for $x\cdot z$:
$(x\cdot z)=\bigl((x\cdot z)\cdot(x_{1}\cdot z_{1})\bigr)\vee\bigl((x\cdot z)\cdot(x_{2}\cdot z_{2})\bigr).$
Here, the method for eliminating the two terms also differs from what was done in [4]. We can obtain p1 and p2 in a similar way. We prove p1:
$x_{1}\cdot z_{1}$ $=((x\cdot x_{1})\vee(c\cdot x_{1}))\cdot z_{1}$ by p1 for $x$
$=(x\cdot x_{1}\cdot z_{1})\vee(c\cdot x_{1}\cdot z_{1}).$ by Mod2
By p1 for $z$ followed by Mod2 on each term of the supremum, the last term is equal to
$(x\cdot x_{1}\cdot z\cdot z_{1})\vee(x\cdot x_{1}\cdot c\cdot z_{1})\vee(c\cdot x_{1}\cdot z\cdot z_{1})\vee(x_{1}\cdot c\cdot z_{1}).$
Due to the observed fact (that $z\cdot x_{1}\cdot z_{1}=x_{1}\cdot z\cdot z_{1}$) and the fact that $c$ is central, we can rewrite this term as
$(x\cdot z\cdot x_{1}\cdot z_{1})\vee(x\cdot c\cdot x_{1}\cdot z_{1})\vee(x_{1}\cdot z\cdot c\cdot z_{1})\vee(c\cdot x_{1}\cdot z_{1}).$ (10)
Note that $(x\cdot c\cdot x_{1}\cdot z_{1})=x\cdot(c\cdot x_{1}\cdot z_{1})\leq(c\cdot x_{1}\cdot z_{1})$, so we can eliminate the second term of the supremum. To eliminate the third term, we can rewrite it as:
$(x_{1}\cdot z\cdot c\cdot z_{1})$ $=(x_{1}\cdot z_{1}\cdot z_{2}\cdot z_{1})$ By prod for $z$
$=(x_{1}\cdot z_{2}\cdot z_{1})$
$=(x_{1}\cdot z\cdot c)$ by prod for $z$
$=(c\cdot x_{1}\cdot z).$ $c$ is central
Let’s see that $(c\cdot x_{1}\cdot z)\leq(c\cdot x_{1}\cdot z_{1})$:
$(c\cdot x_{1}\cdot z)\cdot(c\cdot x_{1}\cdot z_{1})$ $=z\cdot c\cdot x_{1}\cdot z_{1}$
$=z_{1}\cdot z_{2}\cdot x_{1}\cdot z_{1}$ by prod for $z$
$=x_{1}\cdot z_{1}\cdot z_{2}$ by Perm
$=x_{1}\cdot z\cdot c$ by prod for $z$
$=c\cdot x_{1}\cdot z.$ $c$ is central
Then Equation (13) becomes:
$(x\cdot z\cdot x_{1}\cdot z_{1})\vee(c\cdot x_{1}\cdot z_{1})$
and this is equal to
$((x\cdot z)\cdot(x_{1}\cdot z_{1}))\vee(c\cdot(x_{1}\cdot z_{1})).\qed$
4.3 The factorization theorem
We begin by introducing some notation. Recall that $\ker f$ is always a congruence when $f:A\rightarrow B$ is a homomorphism. We say that two congruences $\theta,\delta\in\operatorname{\mathrm{Con}}(A)$ are complementary factor congruences if $\theta\cap\delta=\mathrm{Id}_{A}$ and their relational compositions $\theta\circ\delta$ and $\delta\circ\theta$ equal $A\times A$. Complementary factor congruences are in 1-1 correspondence to direct product decompositions.
As $\langle\langle\cdot,\cdot\rangle\rangle$ defines an isomorphism, there exist canonical projections $\pi_{j}:A\rightarrow I_{j}$ with $j=1,2$ such that:
$\forall x\in A,x_{1}\in I_{1},x_{2}\in I_{2}\ :\ x=\langle\langle x_{1},x_{2}\rangle\rangle\iff\pi_{1}(x)=x_{1}\text{ and }\pi_{2}(x)=x_{2}.$ (11)
Let’s define, for any congruence $\theta$ on $A$, the set $I_{\theta}\mathrel{\mathop{:}}=\{a\in A:a\mathrel{\theta}c\}$.
Theorem 4.5.
Let $A$ be an RRB and $c\in A$ be a central element. The mappings
$\begin{array}[]{ccc}\langle\theta,\delta\rangle&\stackrel{{\scriptstyle\mathsf{I}}}{{\longmapsto}}&\langle I_{\theta},I_{\delta}\rangle\\ \langle\ker\pi_{2},\ker\pi_{1}\rangle&\stackrel{{\scriptstyle\mathsf{K}}}{{\longmapsfrom}}&\langle I_{1},I_{2}\rangle\end{array}$
are mutually inverse maps defined between pairs of complementary factor congruences of $A$ and the set of pairs of subsemigroups $I_{1},I_{2}$ of $A$ such that $A=I_{1}\times_{c}I_{2}$.
Proof.
The only part of this proof that differs in a nontrivial way from [4] is the verification that Mod1 and Mod2 hold.
The mapping $a\mapsto\langle a/\theta,a/\delta\rangle$ is an isomorphism between $A$ and $A/\theta\times A/\delta$. Under this isomorphism, $I_{\theta}$ corresponds to $\{\langle c^{\prime},a^{\prime\prime}\rangle:a^{\prime\prime}\in A/\delta\}$ and $I_{\delta}$ corresponds to $\{\langle a^{\prime},c^{\prime\prime}\rangle:a^{\prime}\in A/\theta\}$, where $c^{\prime}=c/\theta$ and $c^{\prime\prime}=c/\delta$. From now on, we will identify $I_{\theta}$ and $I_{\delta}$ with their respective isomorphic images and verify the axioms for $I_{\theta}\times_{c}I_{\delta}=A$ in $A/\theta\times A/\delta$. Note that since $c$ is central, $c^{\prime}$ and $c^{\prime\prime}$ are also central.
To verify Mod1, suppose $x=\langle x^{\prime},x^{\prime\prime}\rangle$, $y=\langle y^{\prime},y^{\prime\prime}\rangle$, $x_{1}=\langle c^{\prime},x^{\prime\prime}_{1}\rangle\in I_{\theta}$, and $x_{2}=\langle x^{\prime}_{2},c^{\prime\prime}\rangle\in I_{\delta}$. Notice that $x\cdot c\leq x_{1}\cdot x_{2}$ implies
$x^{\prime}\cdot c^{\prime}$ $\leq c^{\prime}\cdot x_{2}^{\prime}$ $x^{\prime\prime}\cdot c^{\prime\prime}$ $\leq x_{1}^{\prime\prime}\cdot c^{\prime\prime}.$ (12)
From the first inequality in (12), we obtain
$(x^{\prime}\cdot c^{\prime})\cdot(x^{\prime}\cdot x_{2}^{\prime})$ $=x^{\prime}\cdot x_{2}^{\prime}\cdot c^{\prime}$ $c^{\prime}$ is central
$=x^{\prime}\cdot c^{\prime}\cdot x_{2}\cdot c^{\prime}$
$=(x^{\prime}\cdot c^{\prime})\cdot(x_{2}^{\prime}\cdot c^{\prime})$
$=x^{\prime}\cdot c^{\prime}.$
In other words, $x^{\prime}\cdot c^{\prime}\leq x^{\prime}\cdot x_{2}^{\prime}$.
Similarly, from the second inequality in (12), we obtain $x^{\prime\prime}\cdot c^{\prime\prime}\leq x^{\prime\prime}\cdot x_{1}^{\prime\prime}$.
That is:
$x^{\prime}\cdot c^{\prime}$ $\leq x^{\prime}\cdot x_{2}^{\prime}$ $x^{\prime\prime}\cdot c^{\prime\prime}$ $\leq x^{\prime\prime}\cdot x_{1}^{\prime\prime},$ (13)
and therefore, we have $x^{\prime}\cdot c^{\prime}\cdot y^{\prime}\leq x^{\prime}\cdot x_{2}^{\prime}\cdot y^{\prime}$ and $x^{\prime\prime}\cdot c^{\prime\prime}\cdot y^{\prime\prime}\leq x^{\prime\prime}\cdot x_{1}^{\prime\prime}\cdot y^{\prime\prime}$. Applying Lemma 4.1, we obtain:
$(x\cdot x_{1}\vee x\cdot x_{2})\cdot y$ $=\langle(x^{\prime}\cdot c^{\prime}\vee x^{\prime}\cdot x_{2}^{\prime})\cdot y^{\prime},(x^{\prime\prime}\cdot x_{1}^{\prime\prime}\vee x^{\prime\prime}\cdot c^{\prime\prime})\cdot y^{\prime\prime}\rangle$
$=\langle x^{\prime}\cdot x_{2}^{\prime}\cdot y^{\prime},x^{\prime\prime}\cdot x_{1}^{\prime\prime}\cdot y^{\prime\prime}\rangle$
$=\langle x^{\prime}\cdot c^{\prime}\cdot y^{\prime}\vee x^{\prime}\cdot x_{2}^{\prime}\cdot y^{\prime},x^{\prime\prime}\cdot x_{1}^{\prime\prime}\cdot y^{\prime\prime}\vee x^{\prime\prime}\cdot c^{\prime\prime}\cdot y^{\prime\prime}\rangle$
$=x\cdot x_{1}\cdot y\vee x\cdot x_{2}\cdot y.$
Now, to distribute to the left, the equations (12) also imply
$y^{\prime}\cdot x^{\prime}\cdot c^{\prime}$ $=y^{\prime}\cdot x^{\prime}\cdot c^{\prime}\cdot x_{2}^{\prime}\cdot c$
$=y^{\prime}\cdot x^{\prime}\cdot x_{2}^{\prime}\cdot c$
$=c\cdot y^{\prime}\cdot x^{\prime}\cdot x_{2}^{\prime}.$
This gives us $y^{\prime}\cdot x^{\prime}\cdot c^{\prime}\leq y^{\prime}\cdot x^{\prime}\cdot x_{2}^{\prime}$ and, analogously, $y^{\prime\prime}\cdot x^{\prime\prime}\cdot c^{\prime}\leq y^{\prime\prime}\cdot x^{\prime\prime}\cdot x_{1}^{\prime\prime}$. Then, together with (13), we have
$y\cdot(x\cdot x_{1}\vee x\cdot x_{2})$ $=\langle y^{\prime}\cdot(x^{\prime}\cdot c^{\prime}\vee x^{\prime}\cdot x_{2}^{\prime}),y^{\prime\prime}\cdot(x^{\prime\prime}\cdot x_{1}^{\prime\prime}\vee x^{\prime\prime}\cdot c^{\prime\prime})\rangle$
$=\langle y^{\prime}\cdot x^{\prime}\cdot x_{2}^{\prime},y^{\prime\prime}\cdot x^{\prime\prime}\cdot x_{1}^{\prime\prime}\rangle$
$=\langle y^{\prime}\cdot x^{\prime}\cdot c^{\prime}\vee y^{\prime}\cdot x^{\prime}\cdot x_{2}^{\prime},y^{\prime\prime}\cdot x^{\prime\prime}\cdot x_{1}^{\prime\prime}\vee y^{\prime\prime}\cdot x^{\prime\prime}\cdot c^{\prime\prime}\rangle$
$=y\cdot x\cdot x_{1}\vee y\cdot x\cdot x_{2}.$
We leave Mod2 to the reader. ∎
As in [4], the characterization of direct representations assume a simpler forms when $c$ is an endpoint of the poset. For example,
Theorem 4.6.
Let $A$ be an RRB with identity $1$. Then $A=I_{1}\times_{1}I_{2}$ if and only if $I_{1},I_{2}\leq A$ satisfy:
  1. Perm
    The elements of $I_{1}$ commute with those of $I_{2}$.
  2. Abs
    For all $x_{1},y_{1}\in I_{1}$ and $z_{2}\in I_{2}$, we have: $x_{1}\vee(y_{1}\cdot z_{2})=x_{1}\vee y_{1}$ (and interchanging $I_{1}$ and $I_{2}$).
  3. Onto
    $I_{1}\cdot I_{2}=A.$
Moreover, $I_{1}$ and $I_{2}$ are filters of AA.∎
5 Conclusion
An order-theoretical point of view while studying bands has proven to be very informative, as the previous results show. We have also seen some fruits of a more general, categorical approach to the study of definability of RRBs and other classes of algebras. This approach has also led us to ask some questions which we now present.
In Example 3.17, we focused on the variety of distributive lattices because it is locally finite, while the free (general) lattice on three generators is infinite. As a result, computing $FU\mathbf{C}$ becomes significantly more complicated when considering general lattices. Observe that the connected components of graphs obtained by applying $U$ contain at most two elements. This motivates omitting the distributivity hypothesis:
Question 5.1.
Let $L:=\{\wedge,\vee,0,1,C\}$ and $K$ the class of bounded (non necessarily distributive) lattices, where $C$ is the binary relation defined by “being a pair of complemented elements”. Which graphs are in the class $\mathcal{K}_{R}$?
For a sample we have (in all cases, a two-element connected component is to be added to the indicated graph):
  • A $3$-cycle, by considering $M_{3}$.
  • The complete graph $K_{n}$, by considering $M_{n}$.
  • In general, the complete bipartite graph $K_{n,m}$ is obtained from $\{0\}+(L_{n}\sqcup L_{m})+\{1\}$, where ++ denotes ordered sum, $\sqcup$ denotes disjoint union, and $L_{k}$ denotes the $k$-element linear order. Note that this allows us to obtain the $4$-cycle.
We do not know if it is possible to obtain the $5$-cycle. We do have an example where one of its components is.
A final last remark is that in our seminal example involving RRBs (Corollary 3.13) we are considering plain poset homorphisms as the morphisms of the category $\mathsf{AP}$. We are very curious if there exists a more appropriate notion of morphism that preserves more of the structure of associative posets.
Acknowledgment
We thank Prof. Miguel Campercholi (Universidad Nacional de Córdoba) for suggesting that we study the existence of a left adjoint for the forgetful functor from the category of RRBs to that of associative posets. We also want to thank Nancy Moyano (Centro de Investigación y Estudios de Matemática, CIEM-FaMAF) for support during this project.